### v1

# An ETH-Tight FPT Algorithm for Rejection-Proof Set Packing with Applications to Kidney Exchange

Abstract
We study the parameterized complexity of a recently introduced multi-agent variant of the Kidney Exchange problem. Given a directed graph $G$ and integers $d$ and $k$, the standard problem asks whether $G$ contains a packing of vertex-disjoint cycles, each of length $\leq d$, covering at least $k$ vertices in total. In the multi-agent setting we consider, the vertex set is partitioned over several agents who reject a cycle packing as solution if it can be modified into an alternative packing that covers more of their own vertices. A cycle packing is called rejection-proof if no agent rejects it and the problem asks whether such a packing exists that covers at least $k$ vertices.
We exploit the sunflower lemma on a set packing formulation of the problem to give a kernel for this $\Sigma_{2}^{P}$-complete problem that is polynomial in $k$ for all constant values of $d$. We also provide a $2^{\mathcal{O}(k\log k)}+n^{\mathcal{O}(1)}$ algorithm based on it and show that this FPT algorithm is asymptotically optimal under the ETH. Further, we generalize the problem by including an additional positive integer $c$ in the input that naturally captures how much agents can modify a given cycle packing to reject it. For every constant $c$, the resulting problem simplifies from being $\Sigma_{2}^{P}$-complete to NP-complete. With a single-exponential algorithm for the setting where $c=1$, we show this to be strictly easier under the ETH than when $c=2$. In turn, we show that any $c\geq 2$ yields a problem that is essentially as hard as the original problem with $c$ unbounded. This displays an interesting discrepancy between the classical and parameterized complexity of the problem and gives a good view of what makes it hard.
keywords:
Parameterized complexity, Multi-agent kidney exchange, Kernelization, Set packing
category:
[…]
1 Introduction
The goal of this paper is to analyze the parameterized complexity of algorithmic problems that arise from a recently introduced multi-party variant of kidney exchange problems. Before describing our results, we first present some relevant context.
For patients with severely reduced kidney function, transplantation is often the preferred treatment [Burra07]. As it is possible to live with only one healthy kidney, live kidney donation is employed to reduce waiting times for sparsely available donor organs. In this setting, a willing donor offers one of their kidneys for transplantation to a patient. The patient and donor must be medically compatible for a successful transplantation. When the patient and donor are not compatible, they can opt to participate in a kidney exchange program (KEP).
Such programs aim to link multiple patient-donor pairs so that a donor from one pair is compatible with a patient from another pair. Their goal is to group sets of patient-donor pairs, such that within each group, a cyclic order of compatibility exists. Performing this cycle of transplants means that donors only undergo surgery if their corresponding patient receives a kidney in that exchange. Individual programs each have their own additional restrictions [Biro21]. It is common to restrict the length of transplantation cycles to some constant $d$, due to practical considerations which include doing the associated surgeries simultaneously.
We study these KEPs from an algorithmic perspective. Given a set of participating patient-donor pairs and the medical compatibilities between them, the algorithmic task becomes to find a combination of transplantation cycles that maximizes the total number of patients receiving a healthy kidney. We can model each patient-donor pair as a vertex and the compatibility of the donor of a pair $a$ to the patient of a pair $b$ as a directed edge from $a$ to $b$. We call a cycle of length at most dddd-cycle and use this notation to obtain the following algorithmic graph problem that describes KEPs in one of their most basic forms.
    Kidney Exchange with $d$-cycles (KE-$d$)     Input: A directed graph $G$ and an integer $k$.     Task: Determine whether there is a set $\mathcal{C}$ of pairwise vertex-disjoint $d$-cycles such that $\mathcal{C}$ covers at least $k$ vertices.
This problem has already been studied from the perspective of parameterized algorithms in the literature, with fixed-parameter tractable (FPT) algorithms w.r.t. the input parameter $k$ and the number of neighborhood equivalence classes as examples [Hebert-JohnsonL24, MaitiD22].
Multi-agent KEPs
Extensions of the standard kidney exchange problem have also been studied. In this work, we consider one such extension which we now motivate and introduce. It is known that the fraction of patients to receive a healthy kidney can increase significantly with an increase in the total number of participating pairs [Agarwal19]. This motivates collaborations between multiple programs, from now on referred to as agents. Such collaborations, however, can be hindered by a mismatch between the individual interests of agents, as these need not align with the collective goal of helping as many patients as possible overall. In particular, a socially optimal solution—in which as many patients as possible are treated overall—is not guaranteed to help as many people from each individual agent as that agent could have achieved on their own. This may stop agents from participating in collaborative programs or, when they do participate, move them to withhold information from the collective.
Several ways of incorporating fairness into multi-agent KEPs have been studied [Agarwal19, Ashlagi15, Biro20, Carvalho23, Carvalho16, Hajaj15, Klimentova21]. Recently, Blom, Smeulders, and Spieksma [Blom24] also introduced a framework to capture fairness via a game-theoretical concept of rejections. They translate this into an algorithmic task that asks for a solution that is not rejected by any agent. We study the resulting algorithmic task. To introduce it, we first define the concept of rejections.
Recall the original KE-$d$ problem on directed graphs. To encode multiple agents into the input, we partition the vertex set of the input graph over the agents. In the obtained graph, we refer to a cycle that only contains vertices from a single agent $i$ as an $i$-internal cycle. Like in the original KE-$d$ problem, a solution consists of a packing $\mathcal{C}$ of pairwise vertex-disjoint $d$-cycles. A given such packing $\mathcal{C}$ will be rejected by an agent under the conditions listed in the definition below. See also Figure 1 for an example, adapted from a figure by Blom et al.
Figure 1: A socially optimal packing of cycles is shown on the left and covers six vertices. However, the blue agent rejects this solution as they could modify it to cover more blue vertices as shown on the right. They can do this by rejecting the left-most 33-cycle and including an internal 22-cycle.
Definition 1.1.
Let $G$ be a graph, where $V(G)=V_{1}\cup\ldots\cup V_{p}$ is partitioned over $p$ different agents. Let $d$ be an upper bound on the length of allowed transplantation cycles. Let $\mathcal{C}$ be a set of pairwise vertex-disjoint $d$-cycles in $G$. If there is an agent $i\in[p]$ for which there exist
  • a subset $\mathcal{C}_{\mathrm{rej}}\subseteq\mathcal{C}$, and
  • a set $\mathcal{C}_{\mathrm{int}}$ of $i$-internal $d$-cycles, such that $\mathcal{C}^{\prime}:=(\mathcal{C}\setminus\mathcal{C}_{\mathrm{rej}})\cup\mathcal{C}_{\mathrm{int}}$ is a set of cycles that are still pairwise vertex-disjoint while $\mathcal{C}^{\prime}$ covers more vertices from $V_{i}$ than $\mathcal{C}$ does,
then, agent $i$ is said to reject the packing $\mathcal{C}$. Equivalently, we may say that agent $i$ specifically rejects the cycles in $\mathcal{C}_{\mathrm{rej}}$ and we refer to $\mathcal{C}^{\prime}$ as an alternative packing by which agent $i$ rejects.
This definition captures that the rejecting agent $i$ could modify a given cycle packing $\mathcal{C}$ into another packing $\mathcal{C}^{\prime}$ that benefits more of their own patients by only making changes pertaining to patients within their control. We call a packing that is not rejected by any agent rejection-proof. Then, a guarantee to always find such a solution also serves as a guarantee for agents that participation in the joint program is never detrimental to their personal interests. This led Blom et al. [Blom24] to introduce the following problem, whose parameterized complexity we investigate in this work.
Problem 1.2.
Rejection-Proof Kidney Exchange with $d$-cycles (RPKE-$d$) A directed graph $G$ where $V(G)=V_{1}\cup\ldots\cup V_{p}$ is partitioned among $p$ agents, and an interger $k$. Determine whether there is a rejection-proof set $\mathcal{C}$ of pairwise vertex-disjoint $d$-cycles such that $\mathcal{C}$ covers at least $k$ vertices. For ease of notation, we refer to any packing of pairwise vertex-disjoint $d$-cycles that covers at least $k$ vertices in an RPKE-$d$ instance as a candidate solution. Hence, a packing of $d$-cycles forms a solution to an instance if it is both a candidate solution and rejection-proof.
The above problem takes a leap in complexity compared to the single-agent problem KE-$d$. Whereas KE-$d$ is known to be NP-complete for all finite $d\geq 3$ [Abraham07], Blom et al. have shown RPKE-$d$ to be $\Sigma_{2}^{P}$-complete for $d\geq 3$, even when there are only two participating agents [Blom24, Theorem 1]. In fact, it can be shown that checking whether a given agent rejects a given candidate solution is already NP-hard since this is at least as hard as checking whether a given packing is optimal in the NP-hard setting of only one agent. Partly motivated by this observation, we also consider a more restricted definition of rejections.
For some integer $c\geq 0$, we say that an agent $c$-rejects a packing $\mathcal{C}$ if the rejection criterion from 1.1 can be satisfied with $|\mathcal{C}_{\mathrm{rej}}|\leq c$. We call a packing that no agent $c$-rejects $c$-rejection-proof. Now, we define the problem $c$-Rejection-Proof Kidney Exchange with $d$-cycles ($c$-RPKE-$d$) as the variant of the above RPKE-$d$ problem in which we aim to determine whether the input contains a candidate solution that is $c$-rejection-proof.
The integer $c$ in this definition indicates some level of complexity that a rejection can have. Additionally, it provides a way to interpolate between the problems KE-$d$ ($c=0$) and RPKE-$d$ ($c=\infty$). More importantly, for any constant $c$, we observe that the task of determining whether a given agent $c$-rejects a given packing $\mathcal{C}$ of $d$-cycles is polynomial-time solvable when $d$ is also constant. This is witnessed by the fact that, if an agent $c$-rejects the packing $\mathcal{C}$, an alternative packing can always be obtained by rejecting a constant number of cycles and including a constant number of internal cycles. It follows that $c$-RPKE-$d$ is contained in NP when $c$ and $d$ are constants, as opposed to the $\Sigma_{2}^{P}$-complete RPKE-$d$ problem that does not limit the number of cycles that agents are allowed to reject.
Our contribution
We study the complexity of RPKE-$d$ and $c$-RPKE-$d$ parameterized by $k$, the number of vertices to be covered, and we present both positive and negative results. We prove our positive results on related set packing problems that generalize the two graph problems. This allows us to use some of the tools developed for such problems.
We give an algorithm that, for constant values of $d$, solves RPKE-$d$ in $2^{\mathcal{O}(k\log k)}+n^{\mathcal{O}(1)}$ time on $n$-vertex graphs. It consists of two parts, the first of which is a kernelization algorithm. It takes an arbitrary instance as input and outputs an equivalent instance of size $\mathcal{O}\left((k^{d+1})^{d}\right)$ in polynomial time, via a careful agent-aware application of the sunflower lemma [Erdos60] to eliminate transplantation cycles. In set packing problems, removing sets during preprocessing typically comes at the risk of removing sets that are required to make an optimal solution, thereby possibly transforming YES-instances into NO-instances. The sunflower lemma provides a tool to circumvent this. Our rejection-proof problem settings introduce an additional type of behavior that needs care: transplantation cycles that do not belong to optimal solutions may still play a role in rejections. Removing the wrong cycles may cause an agent not to reject a certain candidate solution that it otherwise would have. We present a two-step reduction approach guaranteed to preserve the answer. The second part of the algorithm consists of brute-force on the reduced instance.
We give a reduction from Subgraph Isomorphism to show that, under the Exponential Time Hypothesis (ETH), the algorithm is asymptotically optimal in terms of $n$ and $k$. We even show the stronger bound (independent of $k$) that a $2^{o(n\log n)}$ time algorithm for RPKE-$d$ or $c$-RPKE-$d$ would violate the ETH for any $c\geq 2$ and $d\geq 3$. We exploit the mechanism of rejections to facilitate our reduction from the Subgraph Isomorphism problem, for which such a superexponential ETH-based lower bound is known [CyganFGKMPS16]. This gives the rejection-proof kidney exchange problems a similar complexity status to other problems in the literature, such as Metric Dimension and Geodetic Set parameterized by vertex cover number, for which a brute-force algorithm on a polynomial kernel is also ETH-optimal [FoucaudGK0IST25].
To complement the ETH-hardness of $c$-RPKE-$d$ with $c\geq 2$, we give a $3^{n}\cdot n^{\mathcal{O}(1)}$ time algorithm for the problem when $c=1$. This allows us to see surprising jumps in the algorithmic complexity of $c$-RPKE-$d$ for varying values of $c$. While no big difference in complexity is obtained between $c=1$ and $c=0$ (in which case the problem can be solved in $2^{n}\cdot n^{\mathcal{O}(1)}$ time with a standard application of dynamic programming over subsets [CyganFGKMPS16, Chapter 6]), our results show that the problem does become computationally harder from a parameterized perspective when $c=2$ (assuming the ETH). We also see that no more jumps in parameterized complexity appear for larger values of $c$. The surprising aspect here is that this one discrete jump in the parameterized complexity of $c$-RPKE-$d$ does not coincide with the jump in the classical complexity of the problem between being contained in NP vs. not being contained in NP: we have seen that, for constant $d\geq 3$, the $c$-RPKE-$d$ problem is contained in NP for every constant value of $c$, while the RPKE-$d$ problem is $\Sigma_{2}^{P}$-complete.
Another jump in complexity is observed when looking at the number of vertices that are not covered by a solution. When asked to find a solution that covers all vertices in the input, the problem again collapses to the NP-complete packing problem KE-$d$. However, our ETH-hardness proof shows that finding a solution that covers all but one vertex is essentially as difficult as the general problem in terms of running time, assuming the ETH. Combined, our results yield a good understanding of what makes rejection-proof kidney exchange hard.
Organization
The remainder of this work is organized as follows. We start with some preliminaries in section 2. Next, we give a polynomial kernel and an FPT algorithm for a set packing generalization of RPKE-$d$ in LABEL:sec:kernel, and we present a matching ETH-lower bound in LABEL:sec:eth-lower-bound. In LABEL:sec:1-rejection-algo, we give a single-exponential algorithm under $1$-rejections and we conclude in LABEL:sec:conclusion. The proofs of statements marked $(\bigstar)$ are deferred to the appendix.
2 Preliminaries
All graphs considered are directed, unless stated otherwise. We use standard notation for graphs and concepts from parameterized complexity; see the textbook by Cygan et al. [CyganFKLMPPS15]. We write $\langle v_{1},\ldots,v_{\ell}\rangle$ to denote the directed cycle with edge set $\{(v_{1},v_{2}),\ldots,(v_{\ell-1},v_{\ell}),(v_{\ell},v_{1})\}$.
A set system is a pair $(U,\mathcal{S})$ where $U$ is a universe of elements and $\mathcal{S}\subseteq 2^{U}$ is a collection of subsets of this universe. For a subset $U^{\prime}\subseteq U$, we use the notation $\mathcal{S}[U^{\prime}]$ to denote the collection of sets from $\mathcal{S}$ that contain only elements of $U^{\prime}$. A hitting set of $\mathcal{S}$ is a subset of $U$ that intersects each set $S\in\mathcal{S}$. We use the following folklore lemma.
Lemma 2.1.
$(\bigstar)$ For a set system $(U,\mathcal{S})$, it can be checked in $2^{|U|}\cdot(|U|+|\mathcal{S}|)^{\mathcal{O}(1)}$ time whether $\mathcal{S}$ contains a collection of pairwise disjoint sets that covers all elements of $U$.
Next, we present the sunflower lemma by Erdős and Rado [Erdos60] and a related definition of sunflowers in set systems. In a set system $(U,\mathcal{S})$, a sunflower $\mathcal{F}$ with $z$ petals and core $Y$ is a collection of sets $S_{1},\ldots,S_{z}\in\mathcal{S}$ such that $S_{i}\cap S_{j}=Y$ for all $i\neq j$. We refer to a set $S_{i}\setminus Y$ as the petal corresponding to set $S_{i}$ and we require all petals to be non-empty. We state a variant of it as proven in the textbook by Flum and Grohe [FlumGrohe12, page 212].
Lemma 2.2 (Sunflower lemma).
Let $(U,\mathcal{S})$ be a set system, such that each set in $\mathcal{S}$ has size at most $d$. If $|\mathcal{S}|>d\cdot d!(z-1)^{d}$, then the set system contains a sunflower with $z$ petals and such a sunflower can be computed in time polynomial in $|U|$, $|\mathcal{S}|$, and $z$.
The following lemma follows from the definitions of sunflowers and hitting sets.
Lemma 2.3.
$(\bigstar)$ Let $(U,\mathcal{S})$ be a set system with sets of size $\leq d$ for which there exists a hitting set of size $\leq h$, for some integers $d,h$. Let $\mathcal{X}\subseteq\mathcal{S}$ be a collection of disjoint sets and let $\mathcal{F}\subseteq\mathcal{S}$ be a sunflower of size d(h1)+2d(h-1)+2 that contains one of the sets $X\in\mathcal{X}$. Then, there is a set $X^{\prime}\in\mathcal{F}\setminus\{X\}$ such that $(\mathcal{X}\setminus{X})\cup\{X^{\prime}\}$ is a collection of pairwise disjoint sets.
Rejection-proof set packing
As mentioned, we present our positive results for problems that generalize the kidney exchange problems defined in the introduction. Intuitively, one may think of translating an RPKE-$d$ instance on a graph $G$ into this generalized setting by constructing a set system $(U,\mathcal{S})$ with $U=V(G)$ and including a set in $\mathcal{S}$ for every vertex set that forms a $d$-cycle in $G$. Related definitions follow analogously, but we briefly reiterate and name them below. To represent $p$ different agents in a set system $(U,\mathcal{S})$, the universe $U=U_{1}\cup\ldots\cup U_{p}$ is partitioned into $p$ sets. A set $S\in\mathcal{S}$ is called $i$-internal if $S\subseteq U_{i}$.
Definition 2.4.
Let $(U,\mathcal{S})$ be a set system such that the universe $U=U_{1}\cup\ldots\cup U_{p}$ is partitioned over $p$ different agents. Let $\mathcal{X}\subseteq\mathcal{S}$ be a packing of pairwise disjoint sets. Suppose there is an agent $i\in[p]$ for which there exist:
  • a subset $\mathcal{X}_{\mathrm{rej}}\subseteq\mathcal{X}$, and
  • a set $\mathcal{X}_{\mathrm{int}}$ of $i$-internal sets such that $\mathcal{X}^{\prime}:=(\mathcal{X}\setminus\mathcal{X}_{\mathrm{rej}})\cup\mathcal{X}_{\mathrm{int}}$ is a collection of sets that are still pairwise disjoint, while $\mathcal{X}^{\prime}$ covers more elements from $U_{i}$ than $\mathcal{X}$ does,
then, agent $i$ is said to reject the packing $\mathcal{X}$. Equivalently, we may say that agent $i$ specifically rejects the sets in $\mathcal{X}_{\mathrm{rej}}$ and we refer to $\mathcal{X}^{\prime}$ as an alternative packing by which agent $i$ rejects.
We call a packing that is not rejected by any agent rejection-proof, and we use this notation to introduce the following generalized problem statement.